The university registrar claims that fewer than 20 percent of the students who enroll at Ivy University graduate in four years. To test this claim, a random sample of 100 students was selected, and 18 were found to have graduated in four years. Calculate the appropriate test statistic to test the hypotheses.
Answer #1

Sample proportion = 18 / 100 = 0.18

Test statistics

z = - p / sqrt ( p ( 1 - p) / n)

= 0.18 - 0.20 / sqrt( 0.20 * 0.80 / 100)

= -0.50